# Education-Oriented Graph Retrieval-Augmented Generation for Learning Path Recommendation

### v1

@@ 6.6. Knowledge Graph Generation
To demonstrate the effectiveness of the KC prerequisite relationship graph generated by our method, we visualized both the original and generated knowledge graphs for the Junyi and MOOCCubeX datasets. The visualization results of the KC prerequisite relationship graphs are shown in Figure 4. It is important to note that since ASSIST09 does not explicitly provide the KC prerequisite relationship graph, it is not included in the analysis. We observe that in the Junyi dataset, the number of relationships in the generated KC prerequisite relationship graph is higher than in the original graph, indicating that our method captures more and more complex prerequisite relationships between KCs. In the MOOCCubeX dataset, the number of KCs in the generated KC prerequisite relationship graph is greater than that in the original graph, enabling the inclusion of a more comprehensive set of KCs. This helps avoid the issue of insufficient KCs in the graph, which could otherwise limit the applicability of the model. Notably, the number of relationships in MOOCCubeX indicates that each KC is connected to every other KC, which is evidently incorrect for the original graph.
Furthermore, we compared the performance of models using the original KC prerequisite relationship graph and the graph generated by our method on the Junyi dataset. Please note that similarity relationship graphs are uniformly generated by DLELP. From Figure 5, we find that the performance of the model using the generated KC prerequisite relationship graph is comparable to that using the original KC prerequisite relationship graph. This demonstrates that, in addition to automatically generating the prerequisite relationship graph, our method can produce KC prerequisite relationship graphs that perform at a similar level to the original graph, which was manually constructed by experts based on domain knowledge and a deep understanding of the relationships between KCs.
@@ 6.7. Case Study
Figure 7 shows the explanation outputs of the KCs by the LLM (gpt-4o) before and after using TextGrad. When explaining the knowledge concepts ”logical address” and ”virtual address” using the LLM alone, confusion arises, as these two KCs are easily mixed up. The LLM would incorrectly explain that the ”logical address” is the same as the ”virtual address” . After optimizing the explanation process using the TextGrad method, it is clear that the LLM can effectively distinguish between the two KCs.
@@ 7. Conclusion
In this paper, we proposed a DLELP method to achieve a more effective learning path by considering dual KC structures prerequisite relationships and similarity relationships. To be specific, we proposed a knowledge concept structure graph generation module that adaptively generates graphs for knowledge concepts based on diverse educational datasets, enhancing the generalizability of existing learning path recommendation methods. Then, we developed a Discrimination Learning-driven Reinforcement Learning (DLRL) framework that alleviates the blocked phenomenon, further improving the performance of learning path recommendations. Finally, we conducted extensive experiments on three benchmark datasets, demonstrating that our method not only achieves state-of-the-art performance but also provides reasons for the recommended learning paths.

### v2

@@ 5.6. Knowledge Graph Generation
To demonstrate the effectiveness of the KC prerequisite relationship graph generated by KnowLP, we visualized both the original and generated KC structure graphs for the Junyi and MCX datasets. The visualization results of the KC prerequisite relationship graphs are shown in Figure 3. It is important to note that since ASS09 does not explicitly provide the KC prerequisite relationship graph, it is not included in the analysis. We observe that in the Junyi, the number of relationships in the generated KC prerequisite relationship graph is higher than in the original graph, indicating that KnowLP captures more complex prerequisite relationships between KCs. In the MCX, the original knowledge structure graph, as shown in Figure 3 (c), covers only a small portion of the 443 KCs in the dataset. Specifically, it includes 1,035 relationships among just 46 KCs, forming a fully connected graph in which every pair of KCs is linked. However, the relations are largely homogeneous, offering little value for learning path generation due to the lack of meaningful prerequisite differentiation. In contrast, the graph generated by KnowLP incorporates all KCs and maintains a more reasonable number of relationships.
These observations collectively demonstrate that our method generates more informative and structurally sound knowledge concept prerequisite graphs, addressing key limitations in the original datasets, such as missing or incorrect relationships and incomplete concept coverage. Moreover, none of the datasets originally contain similarity relationships, further highlighting the value of our approach in enriching the KC structure to support more effective LPR.
@@ 5.7. Case Study
To validate the effectiveness of the innovative components of our method, we conducted an ablation study. Specifically, we removed the similarity agent module from our method to assess its contribution to the overall performance. As shown in Table LABEL:tab:Ablation_Experiment, when the similarity agent is removed from our method, there is a noticeable performance drop across all three datasets. This highlights the necessity and value of considering the similarity relationships between knowledge concepts in learning path recommendations. Moreover, with the similarity agent included, our model achieves a substantial performance improvement at step=20. This further confirms that similarity relationships play a critical role in unlocking the performance potential of longer learning paths.
@@ 6. Conclusion
In this paper, we proposed a KnowLP method to achieve a more effective learning path by considering dual KC structures prerequisite relationships and similarity relationships. To be specific, we proposed an EDU-GraphRAG module that adaptively generates graphs for knowledge concepts based on diverse educational datasets, enhancing the generalizability of existing learning path recommendation methods. Then, we developed a DLRL module that alleviates the blocked phenomenon, further improving the performance of learning path recommendations. Finally, we conducted extensive experiments on three datasets, demonstrating that our method not only achieves state-of-the-art performance but also generates more effective longer learning paths.